Given that TC = 20,000 + 4 Q + 0.5 Q2

, determine the equations for TFC, TVC, AFC, AVC, AC, 

MC. Confirm that MC = AVC at the minimum point of the AVC curve.


1
Expert's answer
2022-02-23T12:46:23-0500

Given that TC = 20,000 + 4 Q + 0.5 Q2

Equations for:


TFC=20,000; These are costs that do not vary with unit changes.


TVC=4Q+0.5Q2; These costs vary with changes in units of capital.


"AFC=\frac{20000}{Q}"


"AVC=\frac{4Q+0.5Q ^2}{Q}=4+0.5Q"


AC="\frac{20000+4Q+0.5Q ^2}{Q}=\frac{20000}{Q}+4+0.5Q"


MC=4+Q; differentiate total cost function.


Confirm that MC = AVC at the minimum point of the AVC curve.

MC = AVC; 4+Q=4+Q; so, MC=AVC